OpenStudy (anonymous):

Estimate the tip on a $19.65 check using a tip rate of 15%.

OpenStudy (anonymous):

What is 15% of $19.65? Usually when you see the word "of" it indicates multiplication. You do: $19.65 x 0.15 = 2.9475 <-- This is the 15% out of the total $19.65 and is how much you paid for tips.

OpenStudy (anonymous):

You muliply 0.15 , not 15 because we are dealing with percentage. 100% = 100/100= 1.00 15% = 15/100 = 0.15

OpenStudy (anonymous):

multiply*